### v1

# POWSM: A Phonetic Open Whisper-Style Speech Foundation Model

Abstract
Recent advances in spoken language processing have led to substantial progress in phonetic tasks such as automatic speech recognition (ASR), phone recognition (PR), grapheme-to-phoneme conversion (G2P), and phoneme-to-grapheme conversion (P2G). Despite their conceptual similarity, these tasks have largely been studied in isolation, each relying on task-specific architectures and datasets. In this paper, we introduce POWSM (Phonetic Open Whisper-style Speech Model), the first unified framework capable of jointly performing multiple phone-related tasks. POWSM enables seamless conversion between audio, text (graphemes), and phones, opening up new possibilities for universal and low-resource speech processing. Our model outperforms or matches specialized PR models of similar size (Wav2Vec2Phoneme and ZIPA) while jointly supporting G2P, P2G, and ASR. Our training data, code [^1] https://github.com/espnet and models [^2] https://huggingface.co/espnet/powsm are released to foster open science.
POWSM: A Phonetic Open Whisper-Style Speech Foundation Model
Chin-Jou Li1, Kalvin Chang2, Shikhar Bharadwaj1, Eunjung Yeo3, Kwanghee Choi3, Jian Zhu4, David Mortensen1, Shinji Watanabe1, 1Carnegie Mellon University, 2University of California, Berkeley, 3University of Texas, Austin, 4University of British Columbia,
1 Introduction
Figure 1: POWSM is the first phonetic foundation model that can perform four phone-related tasks: Phone Recognition (PR), Automatic Speech Recognition (ASR), audio-guided grapheme-to-phoneme conversion (G2P), and audio-guided phoneme-to-grapheme conversion (P2G).
Phones are the smallest units of sound in speech. Unlike graphemes, phones are shared across languages and usually represented using the International Phonetic Alphabet (IPA) (International Phonetic Association, 1999), a unified transcription standard for all languages. By providing a consistent representation of speech across languages, phone-level modeling allows fine-grained analysis and cross-lingual generalization, enabling tasks like atypical speech analysis (e.g., L2 speech Li et al. (2016); Inceoglu et al. (2023) and pathological speech Choi et al. (2025); Li et al. (2025)), endangered language documentation He et al. (2024), code-switched text-to-speech Zhou et al. (2020), and cross-lingual transfer in speech-to-text Pratap et al. (2024); Magoshi et al. (2025).
Four key phone-related tasks underpin phonetic spoken language processing: automatic speech recognition (ASR), phone recognition (PR), grapheme-to-phoneme conversion (G2P), and phoneme-to-grapheme conversion (P2G). ASR learns implicit phonetic representations Belinkov and Glass (2017), while PR offers explicit phone-level supervision. G2P and P2G bridge orthographic and phonetic spaces. Collectively, these tasks interact through shared phonetic representations, each addressing a different aspect of the relationship between audio, phones, phonemes, and graphemes.
Despite their conceptual similarity, these tasks have traditionally been developed in isolation, using task-specific architectures and datasets. Such systems are optimized for specific input-output mappings and cannot be easily extended to other phonetic tasks. This fragmentation has hindered the development of general-purpose models for phonetic processing, necessitating a unified phonetic foundation model that can perform multiple phone-related tasks within a single, general framework for speech processing.
To bridge this gap, we propose POWSM, a phonetic foundation model capable of performing four core phone-related tasks — PR, ASR, audio-guided G2P, and audio-guided P2G — within one unified architecture (Figure˜1). To construct this framework, we reformulate standard ASR datasets Zhu et al. (2025) into four task-specific formats, allowing the model to learn consistent mappings across audio, phoneme, and grapheme representations. In addition, POWSM adopts an attention-based encoder-decoder (AED) architecture, following the design of large-scale speech foundation models such as Whisper (Radford et al., 2023) and OWSM (Peng et al., 2023).
Empirically, POWSM outperforms previous PR models on both in-domain data and out-of-domain languages, achieves low-resource ASR performance comparable to web-scale multilingual foundation models, and can act as speech-grounded P2G and G2P across more than 70 languages. POWSM offers a new unified paradigm for phone-level modeling, paving the way for inclusive and globally accessible speech technologies that transcend language boundaries and resource disparities.
To summarize, our main contributions are:
  • We provide POWSM, a large-scale foundation model that achieves state-of-the-art PR performance, and is capable of performing multiple fundamental phone-related tasks. Our model enables seamless conversion between speech, text (graphemes/orthography), and phones.
  • We thoroughly analyze POWSM to understand the interaction between multiple tasks, architecture components, and losses.
  • We fully open-source all our data preparation and evaluation scripts, model checkpoints and code to foster open science.
2 Related Work
Speech foundation models
Recent speech foundation models such as Whisper (Radford et al., 2023) and OWSM (Peng et al., 2023, 2024) have driven progress in large-scale multilingual ASR and speech translation, but they do not explicitly address phoneme recognition or articulatory-level supervision. Subsequent work (Yusuyin et al., 2025; Fu et al., 2025) showed that incorporating phoneme-level objectives improves ASR for low-resource and long-tailed settings, while outputting phonemes as an intermediate benefited speech translation (Gállego et al., 2025). POWSM extends this line of work by being the first open foundation model jointly trained on phone recognition and related tasks, integrating multilinguality, phonetic supervision, and multi-task scalability within one framework.
Phone recognition
Prior work in multilingual phone recognition can broadly be categorized into (1) language-specific models (Gao et al., 2021) that rely on explicit phoneme (Xu et al., 2022) or allophone inventories (Li et al., 2020) and (2) language-agnostic approaches that aim to generalize across languages without such resources (Taguchi et al., 2023; Glocker et al., 2023; Li et al., 2021; Zhu et al., 2025). POWSM follows the latter paradigm as a fully data-driven multilingual model that learns phone representations without predefined phoneme mappings.
WhisperPPT (Samir et al., 2025) improved Whisper (Radford et al., 2023)’s performance through data cleaning but remained limited in data coverage and task diversity. However, Whisper is trained on a closed corpus and could display harmful biases for PR which cannot be fully removed by fine-tuning. POWSM is trained from scratch on open datasets.
ZIPA (Zhu et al., 2025) scaled PR to 17,000+ hours of data and 88 languages using a Zipformer Yao et al. (2024) encoder and noisy-student training on 4,000+ languages, achieving state-of-the-art results. To construct its training corpus, ZIPA employed a G2P system to convert large-scale ASR transcriptions into phoneme sequences, effectively repurposing ASR datasets for PR. Building on this idea, POWSM leverages both the grapheme and the G2P-generated phoneme transcriptions, reformulating them into four task-specific forms: ASR, PR, G2P, and P2G.
G2P & P2G POWSM is the first model capable of both audio-guided G2P and audio-guided P2G. G2P conversion, sometimes called phonemization in the text-to-speech literature, can be accomplished with pronunciation dictionaries (Rudnicky, 1993), rules (Mortensen et al., 2018), WFSTs (Black and Lenzo, 2001), or seq2seq neural methods to choose between different pronunciations of a word in context (Zhu et al., 2022). Text-based G2P, however, still cannot handle phonetic variation, enforcing a one-to-one mapping between orthography and transcription. In contrast, audio-guided G2P can learn to map the different acoustic realizations of a phoneme across varieties of a language to a phone representation Route et al. (2019). In particular, Mak et al. (2025) observed a performance improvement in using audio-guided G2P versus text-based G2P alone for Cantonese. Gao et al. (2024) similarly showed that joint learning of G2P, phone recognition, and forced alignment outperform a G2P teacher model. Similarly, Sun and Richmond (2024) jointly learned G2P and TTS. Compared to G2P, P2G conversion is less studied, with Lauc (2024) training a seq2seq model on 19 million language-grapheme-phoneme triplets.
3 Methodology
3.1 Data preparation
We use IPAPack++ Zhu et al. (2025) for training. It is an open source corpus of roughly 17,000 hours of multilingual speech with paired orthographic and phonemic transcriptions. We will release all data processing scripts to make POWSM fully reproducible.
G2P-generated transcriptions have been manually inspected and cleaned. Following Samir et al. (2025), we remove Interlingua and 10 noisy FLEURS languages. Utterances longer than 300 phones are filtered out. IPA sequences are normalized to Unicode NFD (Canonical Decomposition); English G2P sequences are further refined with rule-based corrections to fix voice-onset time issues (see Appendix § A.1).
[…]
3.2 Multitask data format
Our model is trained on four tasks: PR, ASR, and audio-guided G2P and P2G. Each utterance is used once per task, with task-specific formatting as illustrated in Figure˜1, including a text prompt, language token, task token, and target output. We leave the text prompt blank (token <na>) for PR and ASR, and provide graphemes and phones as prompts for G2P and P2G.
3.3 Training details
POWSM adopts an attention-based encoder-decoder (AED) architecture, which flexibly models output sequences and allows the integration of additional tasks. Specifically, we follow the OWSM v3.1 architecture Peng et al. (2024), which employs an E-Branchformer encoder and a Transformer decoder, consistent with the general encoder-decoder structure of Whisper Radford et al. (2023). The model is trained from scratch using ESPnet Watanabe et al. (2018) with a hybrid CTC/attention loss Watanabe et al. (2017), where we set the ratio $\alpha_{\text{ctc}}$ to 0.3:
$\mathcal{L}=\alpha_{\text{ctc}}\mathcal{L}_{\text{ctc}}+(1-\alpha_{\text{ctc}})\mathcal{L}_{\text{attention}}.$ (1)
[…]
4 Experimental Setup
Evaluation metric
We report Phonetic Feature Error Rate (PFER), an edit distance using articulatory features from PanPhon (Mortensen et al., 2016), averaged over the number of phones and computed as in Equation 2 for PR. Each feature contributes $\frac{1}{24}$ distance unit, while insertion and deletion cost 1 unit. The edit distance $D$ grows linearly with the sequence length and has no upper bound.
$\text{PFER}=\frac{1}{\text{\#phone}}\sum_{i=1}^{N}D(\text{feat}(\text{hyp}_{i}),\text{feat}(\text{ref}_{i}))$ (2)
Unlike Phone Error Rate (PER), which considers only exact phone matches, or Phone Token Error Rate (PTER), which treats diacritics and modifiers as separate tokens, PFER computes the edit distance in terms of articulatory features—interpretable subphone attributes (e.g. voicing)—capturing phonetic similarity in a fine-grained fashion. Previous studies Taguchi et al. (2023); Zhu et al. (2025) define PFER as the mean articulatory feature edit distance over the evaluation set. In contrast, we normalize it by the number of phones in the reference transcription to measure the proportion of feature errors per phone.
Decoding hyperparameters
We use a CTC weight (denoted as ctc) of 0.3 and a beam size (denoted as beam) of 3 during decoding for all reported numbers unless specified. Further details on the choice of hyperparameters are discussed in § 6.1.
Evaluation datasets
For unseen languages, we evaluate on three datasets: DoReCo Paschen et al. (2020), VoxAngeles Chodroff et al. (2024), and Tusom2021 Mortensen et al. (2021). DoReCo is a dataset of 50+ languages (with broad transcriptions) intended for documentation of small or endangered languages; we use a 45-language subset. VoxAngeles (Chodroff et al., 2024) is a postprocessed version of the UCLA Phonetics Lab Archive (Ladefoged et al., 2009) containing 95 languages. Tusom is a low-data Tangkhulic language of India not included in the training data. Tusom2021 consists of narrow phonetic transcriptions (unlike the broad transcriptions from G2P on which POWSM was trained) of individual Tusom words. We removed the tones.
[…]
We then evaluated our model on in-domain data from IPAPack++, the dataset seen during training. We followed Zhu et al. (2025) in using LibriSpeech for English, AISHELL for Mandarin, and MLS for European languages, and additionally evaluated on IISc-MILE Tamil A et al. (2022) for Tamil and KSC Khassanov et al. (2021) for Kazakh. For ASR and P2G, we evaluate with FLEURS.
See Table 1 for more details about our evaluation datasets.
PR (In-domain)
eng deu nld fra ita spa
10.58 14.27 12.76 10.07 5.27 10.00
por pol tam kaz cmn
3.74 2.14 16.58 7.07 10.02
PR (Out-of-domain: Unseen languages)
DoReCo VoxA. Tusom.
19.18 1.58 1.16
PR (Out-of-domain: Language variation)
Buckeye DRC-SE L2-ARC EpaDB SO762
7.88 0.77 3.66 2.74 2.32
ASR (FLEURS)
afr orm aze pan tgk mkd
0.66 0.13 2.37 1.48 1.96 2.45
bos slv
2.45 1.76
Table 1: Duration of the test sets for different tasks (in hours). Abbreviated datasets (in order): VoxAngeles, Tusom2021, DoReCo South-England, L2-ARCTIC, SpeechOcean762.
Baselines
We evaluate all PR baselines without further training with IPAPack++. See Appendix § A.2 for more details about training data and language coverage. Allosaurus Li et al. (2020, 2021) uses a phone-level CTC to train a language-agnostic model and applies language-specific allophone-to-phoneme mappings. Wav2Vec2Phoneme (Xu et al., 2022), MultIPA Taguchi et al. (2023) and Allophant Glocker et al. (2023) fine-tune XLS-R (Babu et al., 2022) with different objectives: Wav2Vec2Phoneme maps unseen phonemes using articulatory features, MultIPA leverages high-quality G2P data from seven languages, while Allophant decomposes phones into articulatory features and applies CTC losses for each. ZIPA Zhu et al. (2025) trains ZipFormer (Yao et al., 2024) from scratch on IPAPack++ using CR-CTC and also provides a variant trained with additional pseudo-labeled data (“ZIPA-CR-NS-Large”).
For ASR, we compare POWSM with two series of models: OWSM Peng et al. (2025) and OWLS Chen et al. (2025). We select OWSM-CTC v4 because it is the best-performing model in the series, featuring an encoder-CTC architecture that supports ASR, ST, and LID. For OWLS, we include models with comparable parameter sizes.
5 Results
We found that POWSM’s performance on PR and ASR tasks is comparable or superior to competitive baselines.
5.1 Multi-task performance
Results on the in-domain test sets are presented in Table 2 and Table 3. We provide further discussion of G2P and P2G in § 6.2.
POWSM excels at in-domain phone recognition
From Table 2, we see that POWSM achieves the lowest average PFER in phone recognition, due to the strong language modeling capability of the decoder. We hypothesize that our English data cleaning (Appendix § A.1) may have negatively affected the PFER for Germanic languages due to a mismatch between training and test data. Nevertheless, our approach fills this gap by achieving strong performance on other languages, outperforming models trained on larger datasets.
Model Param. eng deu nld fra ita spa por pol tam kaz cmn Avg.
Allosaurus 11M 6.89 17.67 19.19 20.91 19.02 4.82 19.61 21.21 12.01 20.90 15.28 16.14
Allophant 300M 10.26 9.37 18.39 18.83 7.82 17.37 15.44 7.90 19.32
Wav2Vec2Phoneme 300M 7.70 7.89 12.31 17.73 6.10 3.67 11.65 9.57 15.63 15.30 14.66 11.11
MultIPA 300M 15.81 16.28 18.97 20.19 7.20 6.99 15.04 2.63 10.54 17.71 21.10 13.86
ZIPA-CR-Large 300M 1.63 3.32 3.03 3.23 3.24 1.98 4.01 4.33 4.59 2.31 1.25 2.99
ZIPA-CR-NS-Large 300M 1.40 3.17 2.83 2.92 3.22 1.53 3.40 4.31 4.15 1.87 0.90 2.70
POWSM 350M 2.85 3.37 5.14 3.27 1.81 1.21 2.90 1.36 3.56 2.25 1.15 2.62
Table 2: PFER ($\downarrow$) on the in-domain dataset, IPAPack++. Languages not supported by Allophant are left blank. Some languages were not seen by MultiIPA. Bold indicates the best performance.
POWSM is comparable with web-scale ASR models on low-resource languages
We hypothesize that pre-training with phone recognition benefits low-resource ASR Yusuyin et al. (2025). To choose low-resource languages, we selected languages in IPAPack++ with less than 8 hours of speech in FLEURS to serve as the test set. See § A.3 for details on the amount of data used by different models. For a fair comparison with other multilingual ASR baselines without language-specific components, we use the same decoding hyperparameters ctc=0.0, beam=1.
As shown in Table 3, POWSM (POWSM 0.35B, ASR) is often comparable to models of similar size trained on web-scale data for ASR (OWLS 0.5B). Incorporating phones obtained from PR as text prompts (PR-P2G) significantly decreases WER, making it comparable to or even better than these models. When using gold phone labels for P2G (see analysis in § 6.2), POWSM outperforms other ASR models by a large margin in most cases.
Afroasiatic Turkic Indo-Iranian Balto-Slavic
Model afr orm aze pan tgk mkd bos slv
OWLS 0.5B 102.3 89.0 77.7 59.3 60.4 54.2 59.3 58.6
OWLS 1B 95.7 102.4 67.5 50.0 50.7 46.2 50.0 52.8
OWSM-CTC v4 1B 67.5 92.7 71.2 88.7 57.6 51.2 51.3 60.4
POWSM 0.35B, ASR 86.2 125.3 67.7 83.1 62.8 56.0 56.5 64.5
POWSM 0.35B, PR-P2G 68.8 93.0 66.7 72.8 51.0 48.6 56.9 63.9
Table 3: WER ($\downarrow$) of ASR and PR-P2G on low-resource languages. PR-P2G uses phones predicted by PR as text prompts instead of gold phones. Bold indicates the best performance, and underline indicates the second-best.
5.2 POWSM generalizes well to unseen languages
Table 4 reports PFER on datasets with unseen languages and language variation. Results indicate that POWSM achieves strong performance across these datasets, and handles both dialectal and L2 variation effectively. Notably, our method outperforms ZIPA trained on the same data and even exceeds ZIPA trained with extra pseudo-labeled data, achieving the best results on unseen languages while performing three additional tasks. This shows the effectiveness of our multi-task approach. While POWSM lags behind Wav2Vec2Phoneme on socio-phonetic variations, we attribute this to its self-supervised learning with over 60k hours of speech (from wav2vec 2.0 (Baevski et al., 2020)) prior to the supervised learning stage.
Unseen Languages Language Variation
Model Param. DoReCo VoxAngeles Tusom2021 Avg. Buckeye DRC-SE L2-ARC EpaDB SO762 Avg.
Allosaurus 11M 24.71 30.84 42.02 32.52 15.24 25.36 13.39 19.33 21.61 18.99
Allophant 300M 16.05 24.13 11.91 14.38 18.28 16.95
Wav2Vec2Phoneme 300M 17.25 13.88 31.92 21.02 12.50 18.57 9.86 9.90 13.60 12.89
MultIPA 300M 18.28 15.23 30.53 21.35 18.69 23.31 15.52 15.64 21.34 18.90
ZIPA-CR-Large 300M 17.99 16.95 23.68 19.54 12.04 17.89 9.74 17.38 15.58 14.53
ZIPA-CR-NS-Large 300M 16.82 17.14 23.08 19.01 12.05 17.12 9.69 14.63 18.20 14.34
POWSM 350M 17.06 17.11 21.96 18.71 12.63 18.33 11.32 11.86 17.84 14.40
Table 4: PFER ($\downarrow$) on out-of-domain data. “DRC-SE” stands for DoReCo South-England; “L2-ARC” stands for L2-ARCTIC; “SO762” stands for SpeechOcean762. Unseen language datasets include languages not supported by Allophant; therefore, we do not report results for these datasets.
6 Analysis
In this section, we analyze how POWSM works, focusing on the phonetic-aware encoder and task- and language-specific tokens, which are the defining features of the model.
6.1 Behavior of the speech encoder
The CTC encoder prefers fine-grained phones without suprasegmentals
We observed that mixing phones and orthography as encoder targets hindered training, because the same speech input would have different encoder CTC targets for different tasks. Therefore, we used phones as encoder targets, encouraging general representations of sounds to be shared across languages.
[…]
We run small-scale experiments on a 1k-hour subset of the multi-task data (250 hours of speech repeated across four tasks). We use the validation CER of the encoder-CTC output as a proxy for training efficiency. An earlier drop indicates that the encoder is learning a useful alignment early, which improves representations fed into the decoder and accelerates overall convergence. In Figure 2, PanPhon tokenization without suprasegmentals shows the earliest drop, suggesting that alignment with decoder units aids training, while collapsing suprasegmental distinctions for CTC reduces confusion.
Figure 2: Validation CER of encoder-CTC during training. Removing suprasegmentals for CTC accelerates convergence and reduces vocabulary size for encoder.
Increased encoder weights benefit PR on out-of-domain data
As in other encoder-decoder models Gong et al. (2023); Radford et al. (2023), we expect the encoder of POWSM to capture more general acoustic patterns, while the decoder handles language and task-specific output formats. Therefore, we investigate whether emphasizing the encoder more during different stages of model development affects performance. To balance data diversity with inference compute costs, we selected two smaller datasets from each category with distinct characteristics.
As shown in Table 5, higher CTC decoding weights improve PR performance on out-of-domain data but degrade it on in-domain data, as expected. This echoes Zhu et al. (2025)’s finding that RNN-T (Graves and Jaitly, 2014), an encoder-only speech-to-text model with an autoregressive text prediction network, hurts generalization to unseen patterns of phones (phonotactics). We hypothesize that the decoder is performing implicit language modeling and “smooths” phonetic variation, as Zhu et al. (2025) described.
Next, we examine whether focusing more on the CTC loss through training widens this gap in performance between in-domain and out-of-domain data. We find that fine-tuning with a higher CTC loss weight $\alpha_{\text{ctc}}$ after convergence does not improve out-of-domain performance and can even degrade it. Randomly varying $\alpha_{\text{ctc}}$ for each batch also shows no improvement. In contrast, training with a higher $\alpha_{\text{ctc}}$ from the start benefits the out-of-domain distribution, achieving the lowest PFER on unseen languages with greedy decoding, while the PFER on in-domain data is comparatively higher. These results suggest that assigning a higher weight to the encoder during training and inference improves PR, highlighting a common trade-off between in-domain performance and generalization.
In-domain Out-of-domain
Setup ita pol VoxA. Tusom. DRC-SE EpaDB
Decoding
ctc=0.3 1.66 1.36 17.58 33.52 18.21 11.88
ctc=0.7 1.97 1.37 17.92 24.29 18.05 11.82
ctc=0.9 2.05 1.38 19.27 22.94 17.67 11.80
Pre-training / Fine-tuning
$\alpha_{\text{ctc}}$=0.3 1.81 1.60 16.02 22.47 18.59 11.66
$\rightarrow$ Ft, $\alpha_{\text{ctc}}$=0.5 1.94 1.53 17.78 23.72 18.73 11.82
$\alpha_{\text{ctc}}$=0.7 1.96 1.65 15.40 22.10 18.50 11.62
$\rightarrow$ Ft, $\alpha_{\text{ctc}}$=0.5 2.01 1.57 16.21 22.93 18.41 11.47
$\alpha_{\text{ctc}}$=UU(0.1,0.9) 1.95 1.62 19.29 25.11 18.92 11.33
Table 5: PFER ($\downarrow$) for different CTC weight settings. “Ft” denotes fine-tuning for 5 epochs from the checkpoint above. VoxAngeles and Tusom2021 are abbreviated. Pre-training and fine-tuning rows use ctc=0.3. All setups use beam=1.
Task Buckeye Example
ASR Transcription any holidays at all they just kind of ignore
[…]
6.2 Inspecting Task and Language Tokens
Speech-guided G2P preserves phonetic variation; text prompts normalize it
To better understand how POWSM integrates speech and text prompts, we analyze the relative influence of speech and text prompts in its G2P behavior. We vary the G2P conditions from purely speech-based to purely text-based, as shown in Table 6, and evaluate the model on the Buckeye dataset. When only speech is provided, the performance is comparable to the PR setting, which differs only in the task token. Adding both speech and text prompts (the standard G2P setup) leads to degraded performance, with output showing standardized pronunciations. When the model relies solely on the text prompt, performance drops sharply and pronunciations become highly standardized as expected (just as Zhu et al. (2025) reported).
In other words, POWSM G2P responds to speech and text signals to controllably mediate between narrow and broad transcription. In the multi-task setup, this effect may be stronger because the model is trained with G2P, which could bias it toward more standardized forms.
Audio-P2G effectively handles low-resource languages
We compare several P2G setups on the same set of low-resource languages from FLEURS, listed in Table 7. P2G significantly outperforms ASR, suggesting that it effectively leverages the provided phone context. However, since P2G uses gold phone labels, this comparison is not entirely fair. We therefore tested PR followed by P2G (PR-P2G), and found that performance improved for some languages but not for others. Error propagation does not explain this variation in performance, as PFER trends from PR differ from the observed performance drops. Yet the PFER pattern aligns closely with ASR results, suggesting that phonotactic similarity to high-resource languages may play a role.
To test this, we run P2G with the language code set to English and post-process the output to match Cyrillic or Gurmukhi transcriptions with online conversion tools for certain languages. [^6] https://www.lexilogos.com for Macedonian and Tajik; https://punjabi.indiatyping.com for Panjabi. This approach often outperforms ASR and sometimes approaches P2G’s performance, indicating that P2G also relies heavily on speech input. Languages with either comparibly low or high PFER did not benefit from this transliteration approach, possibly because the model already handled them well or had not yet learned them sufficiently. This finding suggests a direction for further investigation in low-resource ASR.
Afroasiatic Turkic Indo-Iranian Balto-Slavic
Task afr orm aze pan tgk mkd bos slv
Best ASR 67.5 89.0 67.5 50.0 50.7 46.2 50.0 52.8
ASR 86.2 125.3 67.7 83.1 62.8 56 56.5 64.5
P2G 55.9 88.0 37.1 52.6 31.8 36.9 32.3 40.3
P2G, lang=<eng> 60.4 99.0 64.2 95.8 74.0 52.6 39.6 53.5
PR-P2G 68.8 93.0 66.7 72.8 51.0 48.6 56.9 63.9
   PFER ($\downarrow$) 9.1 12.9 6.7 7.9 5.7 3.3 6.7 6.9
Table 7: WER ($\downarrow$) of different P2G settings on low-resource languages “Best” stands for lowest WER in Table 3 from ASR models. indicates post-processed languages.
Language token captures phonotactics
The language identification (LID) performance of POWSM on seen languages in FLEURS reaches 92.3% accuracy, as shown in Figure 3. To see if the model implicitly learns phonotactic patterns and associates them with the language token, we evaluate PR on unseen languages by manipulating the language token at inference time. For VoxAngeles and Tusom2021, the three most frequently assigned languages are Bashkir (42.6%, 25.1%), English (30.2%, 67.7%), and Kinyarwanda (14.5%, 2.3%), which are all relatively high-resource languages in IPAPack++. Table 8 shows that assigning the detected language token yields better performance than always using English, while setting the language as unknown performs best. This indicates that the language token influences PR by shifting the output distribution toward the assigned language.
Lang. token Voxangeles Tusom2021 Example
[…]
Table 8: PFER ($\downarrow$) of PR performance with different language token. The detected language in the example is <bak>. Blue for correct, orange for error compared to other examples.
7 Conclusion
We train a fully open-source phonetic speech foundation model POWSM using our scalable multi-task framework. Our model achieves state-of-the-art performance on PR while also supporting ASR across more than 70 languages. Beyond PR and ASR, the model’s ability to perform audio-guided G2P and P2G enables applications that require fine-grained linguistic analysis such as atypical speech assessment. Our analysis reveals that POWSM’s encoder benefits from phoneme-level CTC supervision and stronger encoder weighting, enhancing cross-lingual generalization. Additionally, the model demonstrates interpretable multimodal and language-aware behaviors, effectively mediating between phonetic detail and standardized phonological patterns. To conclude, POWSM not only provides a strong phone recognition foundation model for high-resource languages, but also acts as a versatile resource for unseen languages and socio-phonetic variation.
8 Future Work
POWSM’s current decoder serves as a large phoneme-level phonotactic language model on which linguists could investigate hypotheses about phonetic universals (Chodroff et al., 2024; Chodroff, 2025) and phonotactics (Shim et al., 2024; Pimentel et al., 2020). In the future, we seek to adapt to socio-phonetic variation either through (unsupervised) test-time adaptation (Lin et al., 2022), in-context learning (Roll et al., 2025; Wang et al., 2024), or mechanistic interpretability (Tang et al., 2024). Furthermore, since Shim et al. (2025) found that earlier encoder layers in Whisper preserve more phonetic detail, early exiting may mitigate the decoder’s tendencies to normalize socio-phonetic variation.
Limitations
POWSM has several limitations that we aim to address in future work. First, the model is neither strictly phonemic nor phonetic: its training data consist of cleaned and filtered phonemic transcriptions from multiple languages, which are not fully faithful to the phonetic or phonemic structure of the audio. Although phonemic transcriptions share similarities across languages, adding auxiliary tasks and language tokens may have reinforced language-specific biases. We also currently lack sufficient allophone-level data, which would provide more language-independent information.
Second, the model still favors high-resource languages. Since we include a decoder for language modeling and language tokens, both of which function effectively, the model would inherently bias toward the seen distribution.
Finally, the current AED architecture imposes engineering limitations. Inference is significantly slower than with encoder-only models, and the architecture does not easily support tone modeling, limiting its application to tonal languages.
Ethics Statement
All of our data is ethically sourced, either through permissive licensing or through proper consent. We are aware of the implicit prescriptivism and representational harms (Crawford, 2017) that normalizing socio-phonetic variation in ASR or PR models can create. This may threaten linguistic diversity instead of preserving it. We also acknowledge that accurate modeling of socio-phonetic variation can enable demographic inference, as demographics and phonetic variation are deeply intertwined (Labov, 1963). We stress that uses of POWSM must align with our vision: a future where advances in spoken language processing and NLP do not leave low-resource varieties behind.
The Use of LLMs
We acknowledge the use of large language models (LLMs) to assist with grammar correction and clarity improvements in writing this paper. All conceptual, methodological, and experimental contributions were developed independently by the authors.
Appendix A Appendix
A.1 Refining English G2P
We observed confusion in plosive voice-onset times on unseen languages in preliminary experiments, which is likely from English G2P data. For instance, broad phonemic transcription in English typically uses /b/ to transcribe the /b/ in /bat/, but its voice onset timing is actually voiceless in Mainstream American English and is closer to [p]. To mitigate this, we apply rule-based refinements to English G2P transcriptions, adjusting plosive voicing and aspiration, lateral velarization, and vowel nasalization.
[…]
A.2 Baseline Implementation
We provide the baselines’ training data source, number of languages covered in the data, and links to model checkpoints or repository in Table 9.
Model Training Data Sources Language Coverage Model checkpoint / GitHub
PR baselines
Allosaurus VoxForge, Japanese CSJ, Hkust 12 xinjli/allosaurus
Li et al. (2020, 2021) Tedlium, Switchboard etc
Allophant Common Voice 10.0 34 kgnlp/allophant
Glocker et al. (2023)
Wav2Vec2Phoneme MLS, Common Voice, 40+ facebook/wav2vec2-xlsr-53-espeak-cv-ft
Xu et al. (2022) Babel
MultIPA Common Voice 11.0 7 ctaguchi/wav2vec2-large-xlsr-japlmthufielta-ipa1000-ns
Taguchi et al. (2023)
ZIPA IPAPack++ 88 lingjzhu/zipa
Zhu et al. (2025) MMS ulab v2., VoxLingua-107 (Pseudo-label) anyspeech/zipa-large-crctc-ns-800k
ASR baselines
OWSM-CTC v4 OWSM v3.2, YODAS 100+ espnet/owsm_ctc_v4_1B
Peng et al. (2025)
OWLS OWSM v3.2, YODAS 150 espnet/owls-scaling-laws-for-speech-recognition-and-translation
Chen et al. (2025)
Table 9: Overview of the baselines for our work.
A.3 FLEURS language selection for ASR
We first filter out languages with more than 8 hours of training data in IPAPack++ Zhu et al. (2025), keeping only those that are also present in FLEURS. Then, following the training data amounts reported in Chen et al. (2025), we further identify the 50 lowest-resource languages to exclude any that may have other substantial sources not included in IPAPack++. This process leaves us with nine languages. We finally exclude ell, as it is comparatively higher-resource and because there are already three other Balto-Slavic languages. Note that other models use strictly more data than ours—not only in terms of dataset count but also because IPAPack++ applies additional data-quality filtering. Table 10 lists the amount of ASR training data for baselines.
Afroasiatic Turkic Indo-Iranian Balto-Slavic
Model afr orm aze pan tgk mkd bos slv
POWSM 2.71 5.11 6.89 4.96 6.52 5.14 7.57 7.19
OWSM-CTC v4 5.54 6.50 10.69 8.30 8.03 8.4 9.96 26.00
OWLS
Table 10: Amount of ASR training data for languages included in ASR comparison (in hours), according to Zhu et al. (2025) and Chen et al. (2025).
Figure 3: Confusion matrix of LID on FLEURS.
A.4 Multi-tasking at Different Scales
Multi-tasking may improve performance by tying acoustic signals to well-defined symbolic representations, yet it may distract the model if the relationships are not learned effectively. We train POWSM with different data and model scales to examine how multitask learning interacts with the setup, and use beam=1 during decoding to speed up inference.
Table 11 shows that there is no clear trend regarding whether multitasking benefits PR performance. PR performance degrades when the model has excessive capacity relative to the available data (too little data), or when it is limited by size (too much data).
Further evidence is needed before concluding that phoneme recognition benefits less from scaling, as we currently lack sufficient data and large model capacity to test this thoroughly. Nevertheless, the model demonstrates the ability to multitask, which represents a promising direction for future work.
Data ( khr) Tasks Params. VoxAngeles Tusom2021 L2-Arctic
0.25 1 100M 27.22 32.59 13.50
0.25 4 100M 26.30 30.32 13.40
0.25 1 300M 20.63 25.83 12.88
0.25 4 300M 23.81 25.91 14.14
17 1 100M 17.88 26.68 11.76
17 2 100M 24.69 49.28 11.35
17 4 100M 30.07 61.89 12.37
17 1 300M 17.08 25.20 10.50
17 2 300M 17.17 23.70 10.47
17 4 300M 17.58 33.52 10.54
Table 11: Comparison of PFER ($\downarrow$) on different seting of tasks and data. 1 task refers to PR, 2 tasks refer to PR+ASR, and 4 tasks include PR, ASR, P2G, and G2P.